OpenStudy (anonymous):

Two separate bacteria populations grow each month and are represented by the functions f(x) = 3x and g(x) = 7x + 6. In what month is the f(x) population greater than the g(x) population?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@pooja195

jigglypuff314 (jigglypuff314):

well let's first figure out what month they might be equal that would be f(x) = g(x) or 3x = 7x + 6 can you solve for x from there?

OpenStudy (anonymous):

i don't think so

jigglypuff314 (jigglypuff314):

3x = 7x + 6

OpenStudy (anonymous):

i think it would be after 2 months

jigglypuff314 (jigglypuff314):

oh wait... \[f(x) = 3x ~or~ f(x) = 3^x\]

OpenStudy (anonymous):

the first one

jigglypuff314 (jigglypuff314):

then I don't think this question is right :-/

OpenStudy (anonymous):

it definitely is

OpenStudy (anonymous):

Two separate bacteria populations grow each month and are represented by the functions f(x) = 3x and g(x) = 7x + 6. In what month is the f(x) population greater than the g(x) population? Month 1 Month 2 Month 3 Month 4

jigglypuff314 (jigglypuff314):

I still don't think it's possible... sorry

OpenStudy (anonymous):

well if you had to guess what would it be

jigglypuff314 (jigglypuff314):

@inkyvoyd @imqwerty @ganeshie8 help

imqwerty (imqwerty):

\[f(x)=3x\]\[g(x)=7x+6\]\[g(x)=3x+4x+6\] we know that 3x=f(x)\[g(x)=f(x)+4x+6\] we can clearly see that g(x) will be greater than f(x) what ever the value of f(x) be g(x) will have 4x+6 more add to it so g(x)>f(x)

imqwerty (imqwerty):

oh wait i didn't read the ques well

OpenStudy (anonymous):

so none of the answers are wrong?

imqwerty (imqwerty):

but nvm the ques is wrong if x is the number of week

OpenStudy (anonymous):

number of months

imqwerty (imqwerty):

yeah g(x) will be greater than f(x) no matter what

OpenStudy (anonymous):

well then what should i put just 2 months or something

OpenStudy (error1603):

Never. Even without the +6, 7x will always be greater than 3x, except if x is negative, which populations cannot be.

imqwerty (imqwerty):

u can also try the options to find out the populations liek no where f(x) is greater than g(x)

OpenStudy (anonymous):

im sorry the 3x is supposed to be 3^x..... like an exponent

OpenStudy (error1603):

oooooh

OpenStudy (anonymous):

lol sorry

imqwerty (imqwerty):

just try the options out this is the easiest way not much thinking not much time waste (B

OpenStudy (anonymous):

i think its 2 months

jigglypuff314 (jigglypuff314):

I asked you that ;-;

imqwerty (imqwerty):

thats wrong

OpenStudy (anonymous):

sorry @jigglypuff314

imqwerty (imqwerty):

what did u do?

OpenStudy (anonymous):

idk lol i have no idea

OpenStudy (anonymous):

maybe 4

imqwerty (imqwerty):

(:

imqwerty (imqwerty):

don't guess u shuld be sure about it

OpenStudy (anonymous):

was that right?

jigglypuff314 (jigglypuff314):

lol try plugging in 1 for x into 3^x then into 7x + 6 and compare them so that would be 3^1 = 3 and 7(1) + 6 = 13 13 > 3 so it can't be 1 you can keep on trying it for the rest of the options

imqwerty (imqwerty):

i'll not tell u if thats correct or not

OpenStudy (anonymous):

it is :)

jigglypuff314 (jigglypuff314):

good for you :)

OpenStudy (anonymous):

can you help with more?

jigglypuff314 (jigglypuff314):

ask it as a new question post ^_^

OpenStudy (anonymous):

ok

imqwerty (imqwerty):

(: